OpenStudy (anonymous):

Write the equivalent rates. 6/45=__/15 = __/1 = __/ 1hour 45min

OpenStudy (anonymous):

what does the 1hr and 45 min have to do with this?

OpenStudy (anonymous):

$6/45min = __/15min = __/1hour = __/1hr 45min

OpenStudy (anonymous):

ohhh.. i see. Alright since u have the 6/45 u can divide it by 3 to get how much u get in 15 min. So 45/3 = 15. 6/3 = 2. So in 2$/15 min. From there u can calculate how much you get in an hour knowing that there is 60 min in an hour. So 60/15 = 4. so you must multiply everything by 4. therefore you get 4 x 2 = $8/hr. as for 1hr.45 min. u can make that into 60 min + 45 min = 105 min. Divide 105/15 and you get 7. therefore you multiply 7 to your 2 and u get 7 x 2 = $14/1hr and 45 min.